A certain radio station broadcasts at 105 megahertz. This is the frequency of oscillation of electrons in the transmitting antenna producing the radio wave. Find the period of the electrons' motion.
Answer #1

T=

Given f105 megaHz

f = 105\times 10^{6} $ Hz

T = \frac{1}{105\times 10^{6}}

T = 9.52380952\times 10^{-9} $ seconds

T = 9.52\times ^{-9} $ seconds

T = 9.52 $ nanoseconds

$ period of the electrons motion is 9.52 nanoseconds $